# Tightening constraints on primordial oscillations with latest ACT and SPT data

### v1

@@ Abstract
The oscillation in primordial power spectrum (PPS), a fingerprint of not only a wide class of models of inflation but new physics, is of significant theoretical interest, and can be imprinted on the cosmic microwave background (CMB). In this work, we present constraints on periodic oscillations in the PPS using the latest ACT DR6 and SPT-3G D1 CMB data with the precise measurements at high multipoles beyond the Planck angular resolution and sensitivity. It is found that the combination of SPT and ACT with Planck CMB dataset significantly tightens the upper bound to $A_{\mathrm{log,lin}}\lesssim 0.029$ at 95% C.L., showing no hint for primordial oscillations, where $A_{\mathrm{log,lin}}$ are the amplitudes of logarithmic and linear oscillation in the PPS, respectively. Our work presents state-of-the-art CMB constraints on primordial oscillations, highlighting the power of the ground-based CMB experiments in constraining physics beyond the simplest slow-roll models.
I Introduction
@@
Inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982); Starobinsky (1980), the standard paradigm of the very early universe, predicts a nearly scale-invariant primordial power spectrum (PPS) of scalar perturbations. Recent Planck cosmic microwave background (CMB) data Aghanim et al. (2020a) strongly support the simplest slow-roll inflationary scenarios, and its combination with recent BICEP/Keck data Ade et al. (2021) has placed stringent constraints on the landscape of models Akrami et al. (2020).
@@
However, physics beyond simplest slow-roll potentials Starobinsky (1992); Adams et al. (2001); Wang et al. (2005); Gong (2005); Slosar et al. (2019); Achúcarro et al. (2022) can result in the oscillations in the PPS, which also are consistent with current data, see also models with null energy condition violation Cai and Piao (2021, 2022); Cai et al. (2024). In particular, the periodic oscillating pattern is of significant theoretical interest, motivated by such as Trans-Planckian effects Martin and Brandenberger (2001); Easther et al. (2002); Danielsson (2002); Baumann and McAllister (2015), axion monodromy inflation in string theory Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010), and primordial standard clocks Chen (2012); Chen et al. (2016, 2019). There have been some earlier searches for such oscillations with CMB data Easther et al. (2005); Meerburg et al. (2014); Meerburg and Spergel (2014); Peiris et al. (2013); Easther and Flauger (2014); Aich et al. (2013). Currently, the most stringent constraints on primordial oscillations were reported by the Planck cooperation based on their CMB data Akrami et al. (2020), $A_{\mathrm{log,lin}}\lesssim 0.038$ at 95% C.L., where $A_{\mathrm{log,lin}}$ are the amplitudes of logarithmic and linear oscillation in the PPS respectively, which recently has been complemented by constraints from large-scale structure (LSS) surveys, e.g.Beutler et al. (2019); Ballardini et al. (2023); Mergulhão et al. (2023).
Recently, both Atacama Cosmology Telescope (ACT) Louis et al. (2025); Calabrese et al. (2025) and South Pole Telescope (SPT) Camphuis et al. (2025), the ground-based CMB experiments, have released their new data, offering the most precise measurements of small-scale polarization spectra to date. Their combination with Planck data yields the tightest CMB constraints, which shows no evidence for physics beyond $\Lambda$CDM. These high-precision $-\ell$ polarization data are expected to be powerful for constraining primordial oscillations, given their lower foreground contamination and sharper transfer functions compared to the temperature spectrum Chluba et al. (2015), see Ref. Antony et al. (2024) for a search for primordial oscillations using the previous SPT-3G data.
@@
In this work, we present state-of-the-art constraints on primordial oscillations leveraging the latest ACT DR6 and SPT-3G D1 data. As shown in Fig. 1, we find that the constraints set by the combined SPT+ACT dataset on primordial oscillations are comparable to those from Planck, and the combination of Planck+SPT+ACT dataset significantly tightens the upper bound to $A_{\mathrm{log,lin}}\lesssim 0.029$ at 95% C.L., showing no hint for primordial oscillations.
@@
II Primordial oscillations
Typically, different new physics beyond the simplest slow-roll models can make the PPS manifest different oscillatory patterns. The sudden changes in the inflaton potential, such as bumps and steps Starobinsky (1992); Adams et al. (2001); Gong (2005); Hazra et al. (2010) or sharp turns Ashoorioon et al. (2009); Achucarro et al. (2011); Konieczka et al. (2014), can imprint linear oscillations onto the PPS, while the resonant models Chen et al. (2008) and the axion monodromy model Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010) can lead to logarithmically oscillating signals in the PPS.
@@ II Primordial oscillations
III Datasets and Methods
In our analysis, inspired by Camphuis et al. (2025), we consider the datasets as follows:
  • Planck: The Plik-lite likelihood for Planck PR3 $-\ell$ temperature and polarization data, and the Commander likelihood for $-\ell$ TT and the SimALL likelihood for $-\ell$ EE Aghanim et al. (2020b), as well as the Planck PR4 CMB lensing data Carron et al. (2022).
  • SPT+ACT: The foreground-marginalized SPT-lite [^1] https://github.com/SouthPoleTelescope/spt_candl_data likelihood for SPT-3G D1 Camphuis et al. (2025), combined with the ACT-lite [^2] https://github.com/ACTCollaboration/DR6-ACT-lite likelihood for ACT DR6 Louis et al. (2025); Calabrese et al. (2025), as well as the ACT DR6 CMB lensing data [^3] We use the actspt3g_baseline variant of https://github.com/qujia7/spt_act_likelihood. Madhavacheril et al. (2024); Qu et al. (2024); MacCrann et al. (2024) and SPT-3G Ge et al. (2025); Qu et al. (2025). Here, we impose a Gaussian prior on the optical depth, $\tau_{\mathrm{reio}}=0.054\pm 0.007$ Aghanim et al. (2020a).
@@ III Datasets and Methods
We perform the Markov chain Monte Carlo (MCMC) analysis using Cobaya Torrado and Lewis (2021). The observables are computed using the cosmological Boltzmann code CLASS Blas et al. (2011). The Gelman-Rubin criterion Gelman and Rubin (1992) for all chains is converged to $R-1<0.02$. The flat priors adopted for relevant parameters are presented in Table 1. The best-fit parameters and corresponding $\chi^{2}$ values are obtained using the global optimization method in PROSPECT Holm et al. (2023).
IV Results
@@
In Table 2, we see that compared to the Planck results, SPT+ACT yields a slightly tighter constraint on the logarithmic oscillating amplitude $A_{\mathrm{log}}$, but a higher upper limit on $A_{\mathrm{lin}}$ due to the long posterior tail. The combined Planck+SPT+ACT dataset yields our most stringent constraints, $A_{\mathrm{log}}<0.0286$ and $A_{\mathrm{lin}}<0.0271$ ($95\%$ C.L.).
Figure 3: The 2D posterior distribution of $A_{l}$-$\omega_{l}$ ($68\%$ and $95\%$ confidence levels) for logarithmic (left) and linear (right) oscillating spectra, obtained using Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively.
@@
$-\ell$ TT $-0.43$ $-0.33$ $0.07$ $-0.48$
$-\ell$ EE $0.41$ $0.00$ $-0.27$ $-1.32$
Planck $-\ell$ $-8.04$ $-10.33$ $-2.74$ $-2.20$
SPT-3G D1 $-3.73$ $0.08$ $-0.57$ $-1.25$
ACT DR6 $-2.68$ $-6.79$ $-4.84$ $-1.24$
CMB lensing $0.51$ $-0.00$ $-0.19$ $-0.11$ $0.04$ $-0.39$
DESI BAO $-1.08$ $0.33$ $-0.32$ $0.26$ $0.13$ $0.56$
Pantheon+ SN $0.38$ $0.09$ $-0.04$ $0.06$ $-0.01$ $-0.11$
@@
Table 3: $\Delta\chi^{2}$ values for the best-fit models with logarithmic and linear oscillations, relative to models with the standard PPS, obtained using the Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively. The CMB lensing data used for each dataset is different, as described in Sec. III.
@@
V Conclusion
The search for primordial oscillations is of significant theoretical interest and would open a crucial observational window into the physics of inflation beyond the simplest slow-roll models. In this work, we have reported the constraints on primordial oscillations with the latest ACT DR6 and SPT-3G D1 datasets. The combination of SPT and ACT with Planck dataset yield our most stringent constraints, the upper limits on the oscillating amplitudes of $A_{\mathrm{log}}<0.0286$ and $A_{\mathrm{lin}}<0.0281$ ($95\%$ C.L.).
@@ V Conclusion
It should be mentioned that our constraint on primordial oscillations is based on the $\Lambda$CDM model which is suffering from the Hubble tension Di Valentino et al. (2025). Though our combined Planck+SPT+ACT dataset showed the scalar spectral index is $n_{s}\approx 0.973$, when the early dark energy resolution Karwal and Kamionkowski (2016); Poulin et al. (2019) of Hubble tension is considered $n_{s}$ will be possible to shift towards $n_{s}=1$ Ye and Piao (2020); Ye et al. (2021); Jiang and Piao (2022); Smith et al. (2022); Jiang et al. (2023, 2024); Wang and Piao (2024); Wang et al. (2024); Poulin et al. (2025), see also Di Valentino et al. (2018); Giarè et al. (2023). Thus re-exploring the constraints on primordial oscillations in light of combined Planck+SPT+ACT dataset and the potential resolutions of Hubble tension will be also significant.
@@ Acknowledgements.
Appendix A Full results of MCMC and best-fit parameters
The full results of our MCMC analysis, including the best-fit parameters and $\chi^{2}$ values for models with primordial oscillations, are presented in the Tables 4, 5 and 6. We can see that the inclusion of primordial oscillations has negligible impact on the results of other cosmological parameters.
@@ Appendix A Full results of MCMC and best-fit parameters
Table 4: The mean (best-fit) ±$\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with the standard PPS.
@@
$A_{\mathrm{log}}$ $<0.0387(0.0196)$ $<0.0335(0.0153)$ $<0.0286(0.0145)$
$\omega_{\mathrm{log}}$ unconstrained$(31.46)$ unconstrained$(36.23)$ unconstrained $(31.44)$
$\phi_{\mathrm{log}}$ unconstrained$(4.39)$ unconstrained$(6.13)$ unconstrained$(4.23)$
@@
Table 5: The mean (best-fit) ±$\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with logarithmic oscillations. For upper limits, we report the one-side $95\%$ confidence level ($2\sigma$).
@@
$A_{\mathrm{lin}}$ $<0.0357(0.0136)$ $<0.0453(0.0086)$ $<0.0271(0.0229)$
$\omega_{\mathrm{lin}}$ unconstrained$(10.92)$ 56($56(5.41)^{+40}_{-50}$ unconstrained$(75.23)$
$\phi_{\mathrm{lin}}$ unconstrained$(4.01)$ unconstrained$(6.20)$ unconstrained$(6.28)$
$10^{9}A_{\mathrm{s}}$ 2.113($2.113(2.104)^{+0.027}_{-0.031}$ $2.143(2.159)\pm 0.024$ 2.141($2.141(2.126)^{+0.024}_{-0.028}$
$n_{\mathrm{s}}$ $0.9701(0.9706)\pm 0.0033$ $0.9739(0.9766)\pm 0.0057$ $0.9734(0.9728)\pm 0.0030$
@@
$100\Omega_{\mathrm{b}}h^{2}$ $2.250(2.243)\pm 0.013$ $2.247(2.247)\pm 0.012$ $2.2458(2.246)\pm 0.0093$
$\Omega_{\mathrm{c}}h^{2}$ $0.11784(0.11780)\pm 0.00063$ $0.11816(0.11785)\pm 0.00067$ $0.11803(0.11805)\pm 0.00062$
$\tau_{\mathrm{reio}}$ 0.0594($0.0594(0.0559)^{+0.0064}_{-0.0074}$ $0.0646(0.0693)\pm 0.0060$ 0.0648($0.0648(0.0605)^{+0.0062}_{-0.0073}$
@@
Table 6: The mean (best-fit) ±$\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with linear oscillations. For upper limits, we report the one-side $95\%$ confidence level ($2\sigma$).

### v2

@@ Abstract
The oscillation feature in primordial power spectrum (PPS), a fingerprint of not only a wide class of models of inflation but new physics, is of significant theoretical interest, and can be imprinted on the cosmic microwave background (CMB). In this work, we present constraints on periodic oscillations in the PPS using the latest ACT DR6 and SPT-3G D1 CMB data with the precise measurements at high multipoles beyond the Planck angular resolution and sensitivity. It is found that the combination of SPT and ACT with Planck CMB dataset significantly tightens the upper bound to $A_{\mathrm{log,lin}}\lesssim 0.029$ at $95\%$ C.L., showing no hint for primordial oscillations, where $A_{\mathrm{log,lin}}$ are the amplitudes of logarithmic and linear oscillation in the PPS, respectively. Our work presents state-of-the-art CMB constraints on primordial oscillations, highlighting the power of the ground-based CMB experiments in constraining physics beyond the simplest slow-roll models.
@@
Introduction.— Inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982); Starobinsky (1980), the standard paradigm of the very early universe, predicts a nearly scale-invariant primordial power spectrum (PPS) of scalar perturbations. Recent Planck cosmic microwave background (CMB) data Aghanim et al. (2020a) strongly support the simplest slow-roll inflationary scenarios, and its combination with recent BICEP/Keck data Ade et al. (2021) has placed stringent constraints on the landscape of models Akrami et al. (2020).
@@
However, physics beyond simplest slow-roll potentials Starobinsky (1992); Adams et al. (2001); Wang et al. (2005); Gong (2005); Slosar et al. (2019); Achúcarro et al. (2022) can result in the oscillations in the PPS, which also are consistent with current data, see also models with null energy condition violation Cai and Piao (2021, 2022); Cai et al. (2024). In particular, the periodic oscillation feature is of significant theoretical interest, motivated by such as Trans-Planckian effects Martin and Brandenberger (2001); Easther et al. (2002); Danielsson (2002); Baumann and McAllister (2015), axion monodromy inflation in string theory Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010), and primordial standard clocks Chen (2012); Chen et al. (2016, 2019). There have been some earlier searches for such oscillations with CMB data Easther et al. (2005); Meerburg et al. (2014); Meerburg and Spergel (2014); Peiris et al. (2013); Easther and Flauger (2014); Aich et al. (2013). Currently, the most stringent constraints on primordial oscillations were reported by the Planck cooperation based on their CMB data Akrami et al. (2020), $A_{\mathrm{log,lin}}\lesssim 0.038$ at $95\%$ C.L., where $A_{\mathrm{log,lin}}$ are the amplitudes of logarithmic and linear oscillation in the PPS respectively, which recently has been complemented by constraints from large-scale structure (LSS) surveys, e.g.Beutler et al. (2019); Ballardini et al. (2023); Mergulhão et al. (2023). See also Braglia et al. (2021, 2022a, 2022b, 2023); Petretti et al. (2025) for constraints on other realistic models of primordial features.
[…]
@@
In this work, we present state-of-the-art constraints on primordial oscillations leveraging the latest ACT DR6 and SPT-3G D1 data. As shown in Fig. 1, we find that the constraints set by SPT+ACT on primordial oscillations are comparable to those from Planck, and the combined Planck+SPT+ACT dataset significantly tightens the upper bound to $A_{\mathrm{log,lin}}\lesssim 0.029$ at $95\%$ C.L., showing no hint for primordial oscillations.
@@
Primordial oscillations.— Typically, different new physics beyond the simplest slow-roll models can make the PPS manifest different oscillatory patterns. The sudden changes in the inflaton potential, such as bumps and steps Starobinsky (1992); Adams et al. (2001); Gong (2005); Hazra et al. (2010) or sharp turns Ashoorioon et al. (2009); Achucarro et al. (2011); Konieczka et al. (2014), can imprint linear oscillations onto the PPS, while the resonant models Chen et al. (2008) and the axion monodromy model Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010) can lead to logarithmically oscillating signals in the PPS.
@@
Datasets and Methods.— In our analysis, inspired by Camphuis et al. (2025), we consider three datasets as follows:
  • Planck: The Plik-lite likelihood for Planck PR3 high-$\ell$ temperature and polarization data, and the Commander likelihood for low-$\ell$ TT and the SimALL likelihood for low-$\ell$ $E$ Aghanim et al. (2020b), as well as the Planck PR4 CMB lensing data Carron et al. (2022).
  • SPT+ACT: The foreground-marginalized SPT-lite [^1] https://github.com/SouthPoleTelescope/spt_candl_data likelihood for SPT-3G D1 Camphuis et al. (2025); Balkenhol et al. (2024), combined with the ACT-lite [^2] https://github.com/ACTCollaboration/DR6-ACT-lite likelihood for ACT DR6 Louis et al. (2025); Calabrese et al. (2025), as well as the ACT DR6 CMB lensing data [^3] We use the actspt3g_baseline variant of https://github.com/qujia7/spt_act_likelihood. Madhavacheril et al. (2024); Qu et al. (2024); MacCrann et al. (2024) and SPT-3G Ge et al. (2025); Qu et al. (2025). Here, we impose a Gaussian prior on the optical depth, $\tau_{\mathrm{reio}}=0.054\pm 0.007$ Aghanim et al. (2020a).
@@
We perform the Markov chain Monte Carlo (MCMC) analysis using Cobaya Torrado and Lewis (2021). The observables are computed using the cosmological Boltzmann code CLASS Blas et al. (2011). The Gelman-Rubin criterion Gelman and Rubin (1992) for all chains is converged to R1<0.02R-1<0.02. The flat priors adopted for relevant parameters are presented in Table 1. The best-fit parameters and corresponding $\chi^{2}$ values are obtained using the global optimization method in PROSPECT Holm et al. (2023).
Results.— In Table 2, we see that compared to the Planck results, SPT+ACT yields a slightly tighter constraint on the logarithmic oscillating amplitude $A_{\mathrm{log}}$, but a higher upper limit on $A_{\mathrm{lin}}$ due to the long posterior tail. The combined Planck+SPT+ACT dataset yields our most stringent constraints, $A_{\mathrm{log}}<0.0286$ and $A_{\mathrm{lin}}<0.0267$ ($95\%$ C.L.).
@@
Figure 3: The 2D posterior distribution of $A_{l}$-$\omega_{l}$ ($68\%$ and $95\%$ confidence levels) for logarithmic (upper) and linear (lower) oscillating spectra, obtained using Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively.
@@
low-$\ell$ $T$ 0.43-0.43 0.33-0.33 $0.07$ 0.48-0.48
low-$\ell$ $E$ $0.41$ $0.00$ 0.27-0.27 1.32-1.32
Planck high-$\ell$ 8.04-8.04 10.33-10.33 2.74-2.74 2.20-2.20
SPT-3G D1 3.73-3.73 $0.08$ 0.57-0.57 1.25-1.25
ACT DR6 2.68-2.68 6.79-6.79 4.84-4.84 1.24-1.24
CMB lensing $0.51$ 0.00-0.00 0.19-0.19 0.11-0.11 $0.04$ 0.39-0.39
DESI BAO 1.08-1.08 $0.33$ 0.32-0.32 $0.26$ $0.13$ $0.56$
Pantheon+ SN $0.38$ $0.09$ 0.04-0.04 $0.06$ 0.01-0.01 0.11-0.11
@@
Table 3: $\Delta\chi^{2}$ values for the best-fit models with logarithmic and linear oscillations, relative to models with the standard PPS, obtained using the Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively. Note that the CMB lensing data used for each dataset is different.
@@
Conclusion.— The search for primordial oscillations is of significant theoretical interest and would open a crucial observational window into the physics of inflation beyond the simplest slow-roll models. In this work, we have reported the constraints on primordial oscillations with the latest ACT DR6 and SPT-3G D1 datasets. The combination of SPT and ACT with Planck dataset yield our most stringent constraints, the upper limits on the oscillating amplitudes of $A_{\mathrm{log}}<0.0286$ and $A_{\mathrm{lin}}<0.0267$ ($95\%$ C.L.).
@@
[…]
@@ Acknowledgements.
This work is supported by NSFC, No.12475064, 12075246, National Key Research and Development Program of China, No. 2021YFC2203004, and the Fundamental Research Funds for the Central Universities. We acknowledge the use of high performance computing services provided by the International Centre for Theoretical Physics Asia-Pacific cluster and the Tianhe-2 supercomputer.
Supplemental Material for “Tightening constraints on primordial oscillations with latest ACT and SPT data”
I Full results of MCMC and best-fit parameters
The full results of our MCMC analysis, including the best-fit parameters and $\chi^{2}$ values for models with primordial oscillations, are presented in the Tables 1, 2 and 3. We can see that the inclusion of primordial oscillations has negligible impact on the results of other cosmological parameters.
@@ I Full results of MCMC and best-fit parameters
Table 1: The mean (best-fit) $\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with the standard PPS.
@@
$A_{\mathrm{log}}$ <0.0387(0.0196)<0.0387(0.0196) <0.0335(0.0153)<0.0335(0.0153) <0.0286(0.0145)<0.0286(0.0145)
$\omega_{\mathrm{log}}$ unconstrained(31.46)(31.46) unconstrained(36.23)(36.23) unconstrained $(31.44)$
$\phi_{\mathrm{log}}$ unconstrained(4.39)(4.39) unconstrained(6.13)(6.13) unconstrained(4.23)(4.23)
@@
Table 2: The mean (best-fit) $\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with logarithmic oscillations. For upper limits, we report the one-side $95\%$ confidence level ($2\sigma$).
@@
$A_{\mathrm{lin}}$ <0.0357(0.0136)<0.0357(0.0136) <0.0433(0.0086)<0.0433(0.0086) <0.0267(0.0229)<0.0267(0.0229)
$\omega_{\mathrm{lin}}$ unconstrained(10.92)(10.92) $56(5.41)^{+40}_{-50}$ unconstrained(75.23)(75.23)
$\phi_{\mathrm{lin}}$ unconstrained(4.01)(4.01) unconstrained(6.20)(6.20) unconstrained(6.28)(6.28)
$10^{9}A_{\mathrm{s}}$ $2.113(2.104)^{+0.027}_{-0.031}$ $2.142(2.159)\pm 0.024$ $2.141(2.126)^{+0.024}_{-0.028}$
$n_{\mathrm{s}}$ $0.9701(0.9706)\pm 0.0033$ $0.9742(0.9766)\pm 0.0057$ $0.9734(0.9728)\pm 0.0030$
@@
$100\Omega_{\mathrm{b}}h^{2}$ $250(2.243)\pm 0.013 2.2$$47(2.247)\pm 0.012 2.2$$2457(2.246)\pm 0.0094$
$\Omega_{\mathrm{c}}h^{2}$ $0.11784(0.11780)\pm 0.00063$ $0.11816(0.11785)\pm 0.00067$ $0.11803(0.11805)\pm 0.00061$
$\tau_{\mathrm{reio}}$ $0594(0.0559)^{+0.0064}_{-0.0074} 0.0$$645(0.0693)\pm 0.0059 0.0$$0649(0.0605)^{+0.0062}_{-0.0073}$
@@
Table 3: The mean (best-fit) $\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with linear oscillations. For upper limits, we report the one-side $95\%$ confidence level ($2\sigma$).